A multilane highway has four lanes (two lanes in each direction) and a measured free-flow speed of 55 mi/h. One upgrade is 3% and 2 miles long. Trucks are not permitted on the highway, but there are 5% buses in the directional peak hour volume of 1900 vehicles with a peak hour factor of 0.80. What is the calculated density of the highway, assuming familiar drivers are in the traffic stream?
Answer #1

Given,

Free Flow Speed, FFS = 55 mph

Number of lanes in each direction = 2

Peak Flow, V = 1900 veh/hr

Peak-hour factor = 0.80

No Trucks

Buses = 5 %

3% upgrade and 2 miles long

EB = 2.5

fHV = 1/ (1 + 0.05 (2.5-1)) = 1/1.075 = 0.93

fP = 1.0

Peak Flow Rate, Vp = V / (PHV*n*fHV*fP) = 1900/(0.93*2*0.80*1.0) = 1276.88 ~ 1277 veh/hr/ln

Vp < 1400

S = FFS = 55 mph

Density = Vp/S = (1277) / (55) = 23.22 veh/mi/ln

Hence Level of Service is C

Density of LOS C should lie between 18 – 26 veh/mi/ln